Express the quantity 487.4×10^(−12) s in each unit. (ms, ns, ps, fs)
Answer #1

487.4×10−12s  

1second = 1000milliseconds

1s   = 1000ms

487.4×10−12s   = 487.4×10−12*1000ms

                          = 487.4*10^-9 ms

487.4×10−12s  

1second = 10^9nanoseconds

1s   = 10^9 ns

487.4×10−12s   = 487.4×10−12*10^9ns

                          = 487.4*10^-3 ns

487.4×10−12s  

1second = 10^12picoseconds

1s   = 10^12 ps

487.4×10−12s   = 487.4×10−12*10^12ps

                          = 487.4 ps

487.4×10−12s  

1second = 10^15femtoseconds

1s   = 10^15 fs

487.4×10−12s   = 487.4×10−12*10^15fs

                          = 487.4*10^3 fs